# Intrinsic Mixed-state Quantum Topological Order

### v2

@@ Abstract
Decoherence is a major obstacle to the preparation of topological order in noisy intermediate-scale quantum devices. Here, we show that decoherence can also give rise to new types of topological order. Specifically, we construct concrete examples by proliferating fermionic anyons in the toric code via local quantum channels. The resulting mixed states retain long-range entanglement, which manifests in the nonzero topological entanglement negativity, though the topological quantum memory is destroyed by decoherence. By comparison to the gapless spin liquid in pure states, we show that the identified states represent a novel intrinsic mixed-state quantum topological order, which has no counterpart in pure states. Through the lens of quantum anomalies of 1-form symmetries, we then provide general constructions of intrinsic mixed-state quantum topological order, and reveal the existence of non-bosonic deconfined anyons as another key feature of these novel phases. The extended meaning and characterization of deconfined excitations and their statistics in mixed states are clarified. Moreover, when these deconfined anyons have nontrivial braiding statistics, we prove that the mixed states cannot be prepared via finite-depth local quantum channels from any bipartite separable states. We further demonstrate our construction using the decohered Kitaev honeycomb model and the decohered double semion model. In the latter case, a surprising scenario arises where decoherence gives rise to additional types of deconfined anyons.
@@ I Introduction
As long-range entangled (LRE) quantum matter, topologically ordered phases have attracted extensive attention in the past few decades WEN (1990); Chen et al. (2010); Wen (2017); Savary and Balents (2016); Zhou et al. (2017); Sachdev (2018). Recently, there is a growing number of theoretical proposals Aguado et al. (2008); Verresen et al. (2021); Piroli et al. (2021); Tantivasadakarn et al. (2021, 2022, 2023); Lee et al. (2022a); Bravyi et al. (2022); Lu et al. (2023) as well as experimental evidence Semeghini et al. (2021); Satzinger et al. (2021); Andersen et al. (2023); Xu et al. (2023); Iqbal et al. (2023a, b) showing that topological order (TO) can be prepared in current many-body quantum simulation platforms, such as superconducting qubit arrays, Rydberg atom arrays, and trapped ion systems, etc. A key feature of these noisy intermediate-scale quantum (NISQ) devices is the inevitable presence of decoherence, which renders the quantum state a mixed state Preskill (2018). Substantial progress has been made in diagnosing nontrivial topological phases subject to decoherence Dennis et al. (2002); Bombin et al. (2012); McGinley and Cooper (2020); Deng et al. (2021); Wang et al. (2021); de Groot et al. (2022); Lee et al. (2022b); Zhang et al. (2022); Fan et al. (2023); Bao et al. (2023); Lee et al. (2023); Ma and Wang (2023); Ma et al. (2023); Bardyn et al. (2018); Mao et al. (2023); Su et al. (2023); Wang et al. (2023); Liu and Lieu (2024). Particularly, the decoherence-induced breakdown of topological quantum memory in the toric code model is investigated, and is related to the transition in the mixed-state topological order Fan et al. (2023); Lee et al. (2023). The topological order therein is inherited from the pure-state counterpart, which is resistant to modest decoherence. Above certain critical error rate, the long-range entanglement is destroyed.
@@
We are interested in the following question: Other than destroying the pure-state topological order, can decoherence give rise to novel types of quantum topological order that are intrinsically mixed? The possibility of such intriguing scenario arises from new mechanisms of anyon proliferation provided by decoherence, distinct from anyon condensation in pure states Bais and Slingerland (2009); Kong (2014); Burnell (2018). As a starting point, we explore such possibility in the context of $\mathbb{Z}_{2}$ (toric-code) topological order, which comprises three types of anyon excitations, $e,m,$ and $f=e\times m$, Anderson (1973); Read and Chakraborty (1989); Read and Sachdev (1991); Wen (1991); Anderson (1987); Senthil and Fisher (2000); Moessner and Sondhi (2001); Kitaev (2003). In pure states, either $e$ or $m$ (both being self-bosons) can condense, leading to a topologically trivial Higgs/confined phase Fradkin and Shenker (1979). Correspondingly, proliferation of $e,m$ anyons induced by decoherence also destroys long-range entanglement. In contrast, $f$ anyons are self-fermions, and therefore they cannot condense in pure state; instead, strong fluctuation of $f$ anyons typically leads to a gapless spin liquid, which remains LRE. This motivates us to study the fate of $\mathbb{Z}_{2}$ topological order when $f$ anyons proliferate under decoherence.
@@
Specifically, we study the behavior of the toric code model Kitaev (2003) under local quantum channels that solely create $f$ anyons. The topological quantum memory degrades to classical memory above certain decoherence threshold, aligning with previous studies. However, it turns out that the mixed state still possesses nontrivial quantum topological order even in the absence of quantum memory. To diagnose the mixed-state topological order, we employ the topological entanglement negativity (TEN) Lee and Vidal (2013); Castelnovo (2013); Wen et al. (2016a, b), which is a natural generalization of the topological entanglement entropy (TEE) Kitaev and Preskill (2006); Levin and Wen (2006). TEN has been utilized to probe topological order in thermal equilibrium Hart and Castelnovo (2018); Lu et al. (2020, 2022); Lu and Vijay (2022) or under shallow-depth noise channels Fan et al. (2023), effectively acting as a faithful indicator of topological quantum memory in those studied cases. However, we find that the TEN fails to reflect topological quantum memory in this scenario, and remains unchanged across the transition. Nevertheless, the nonzero TEN still points to the persistence of long-range entanglement, a hallmark of quantum topological order. Moreover, the absence of topological quantum memory indicates that the identified topological order has no pure-state counterpart and, therefore, is termed “intrinsic mixed-state quantum topological order” here.
@@
We then provide further understanding of this peculiar result from the perspective of quantum anomalies. Crucially, the noisy channel proliferating $f$ anyons preserves an anomalous $1$-form symmetry generated by $f$ anyons Gaiotto et al. (2015); Wen (2019); McGreevy (2023). Through the anomalous $1$-form symmetry, we show that the intrinsic mixed-state quantum TO in the decohered toric code supports deconfined fermionic anyons, with a detailed explanation of the extended meaning of deconfinement and fermionic statistics in mixed states. We generally prove that mixed states with anomalous $1$-form symmetries must be LRE. Moreover, we show that when the anomalous $1$-form symmetry is generated by anyons with nontrivial braiding statistics, any two complementary parts of the mixed state are LRE (for arbitrary bipartition), generalizing the conjecture “mixed-state anomaly $\Rightarrow$ multipartite non-separability” in Lessa et al. (2024), which focuses on 00-form symmetries. With this perspective, our construction can be straightforwardly generalized to get other intrinsic mixed-state quantum TO, characterized by deconfined anyons with nontrivial statistics. We give two more examples, the decohered Kitaev honeycomb model and the decohered double semion model. In the latter example, we find more surprising features in mixed-state TO. In addition to peeling off anyons in the original TO, decoherence can even give birth to new types of deconfined anyons, which leads to a non-modular anyon theory in our case.
@@
The rest of the paper is organized as follows. In Section II we construct an intrinsic mixed-state quantum TO by proliferating $f$ anyons in the toric code, and reveal its properties through investigations of information quantities including the coherent information and the TEN. We also discuss similarities and disparities to the gapless spin liquid phase in pure states. In Section III we discuss general aspects of intrinsic mixed-state from the perspective of 1-form symmetry anomalies and deconfined excitations. In Section IV we generalize our construction of mixed-state quantum TO to decohered Kitaev honeycomb model and decohered double semion model. We conclude with a discussion in Section V.
II Decohered Toric code as intrinsic mixed-state quantum TO
@@ II.1 The model
$\mathcal{N}^{x}$ $=\prod_{i}\mathcal{N}_{i}^{x},\mathcal{N}_{i}^{x}[\cdot]\equiv(1-p_{x})\cdot+p_{x}X_{i}\cdot X_{i},$ (1)
$\mathcal{N}^{z}$ $=\prod_{i}\mathcal{N}_{i}^{z},\mathcal{N}_{i}^{z}[\cdot]\equiv(1-p_{z})\cdot+p_{z}Z_{i}\cdot Z_{i}.$
@@
where $\bm{\delta}=(\frac{1}{2},-\frac{1}{2})$ (the lattice constant is taken to be 1) and $0<p_{f}<\frac{1}{2}$ is the error rate. In this way the ground state $\rho_{0}$ is turned into a mixed state, $\rho_{f}=\mathcal{N}^{f}[\rho_{0}]$. As depicted in Fig. 1(b), $\mathcal{N}^{f}$ exclusively creates $f$ anyons. In contrast, certain other types of errors, like the Pauli-Y errors, locally create pairs of $f$ anyons as well, but globally they also produce $e$ and $m$ anyons, resulting in completely different outcomes. In the following sections, we demonstrate that this simple model surprisingly realizes an exotic intrinsic mixed-state quantum topological order through analytical exact investigations of its topological memory and topological entanglement negativity.
@@ II.3 Topological entanglement negativity
Based on the above analysis, it may seem that $\rho_{f}$ closely resembles $\rho_{e}$ and $\rho_{m}$. However, surprisingly, we demonstrate below that even when the quantum memory breaks down for $p_{f}>p_{c}$, $\rho_{f}$ retains LRE with a nonzero TEN, indicating the emergence of a distinct quantum topological order.
@@
$G_{x}\equiv\langle\{A_{v}\}\rangle,G_{z}\equiv\langle\{B_{p}\}\rangle.$ (6)
@@
The dependence of TEN on the boundary size for the second type of bipartition seems a bit puzzling. However, for pure-state topological order, similar behavior of the even/odd dependence on system size also exist. Namely, for $\mathbb{Z}_{2}$ topological order enriched by translation symmetry through weak symmetry breaking, typical examples include the Wen-plaquette model and the Abelian phase of Kitaev honeycomb model Wen (2003); Kitaev (2006). In these cases, it is well known that the ground state degeneracy on a torus also exhibits similar dependence on the system size. Although there is little discussion in literature about how weak symmetry breaking affects entanglement properties, it is straightforward to check that the subleading term of bipartite entanglement entropy on a cylinder also depends on the parity of the boundary size in these models. Based on this observation, we establish a connection between the entanglement properties of $\rho_{f}$ and ground-state $\mathbb{Z}_{2}$ topological order enriched by translation symmetry in Appendix A.4.
@@
II.4 Robustness of the intrinsic mixed-state quantum TO
Although in our construction we need to use specific $2$-qubit channels that looks a bit unconventional, $\rho_{f}$ represents a new type of topologically ordered phase, instead of a fine-tuned exception. We can consider the case when single-qubit phase errors are also present: ${\rho}_{f,e}=\mathcal{N}^{z}[\rho_{f}]$, with error rate $p_{z}$. By mapping to two decoupled RBIMs (see Appendix A.5 for details), we demonstrate that for small $p_{z}$, ${\rho}_{f,e}$ stays in the same phase as $\rho_{f}$, while for $p_{z}>p_{c}\approx 0.109$, the state undergoes another transition to the trivial phase, with no memory and zero TEN.
Figure 2: The phase diagram of the toric code model subject to both two-qubit errors and single-qubit phase errors. Properties of each phase, including topological memory, TEN, and superselection sectors are indicated in the phase diagram.
@@ II.5 Comparison to anyon condensation and gapless spin liquid
The distinction becomes much more significant for $f$ anyons. As fermions, they cannot condense in pure states. Then what happens when we turn on fluctuations of $f$ anyons ($h_{zx}\neq 0$)? Surprisingly, at $h_{z}=h_{x}=0$, this model can be exactly solved via fermionization Chen et al. (2018). As we show in Appendix B, when $h_{xz}$ is sufficiently large ($h_{x,z}=0$), corresponding to large fluctuations of $f$ particles, the system enters a gapless spin liquid phase, where $f$ particles form a $p-$wave superconductor with conic dispersion, similar to the gapless Kitaev spin liquid Kitaev (2006). The intrinsic mixed-state quantum TO proposed in this paper has many common features with the gapless spin liquid. First, they are both obtained by proliferating $f$ anyons in the $\mathbb{Z}_{2}$ TO, and, as a result, they both have no quantum memory. For the gapless spin liquid, it is due to the absence of a spectrum gap, so there is no well-defined topological degeneracy. Second, despite the lack of quantum memory, they are both LRE, and are thus nontrivial phases of matter. In the next section we will uncover a deeper reason of their LRE nature from the perspective of anomalies. However, there are also notable differences between $\rho_{f}$ and the gapless spin liquid phase. Gapless phases typically exhibit critical behaviors, including algebraically decaying correlation functions and subleading logarithmic corner contributions to entanglement entropy/negativity Fradkin and Moore (2006); Casini et al. (2009); Bueno et al. (2015). Contrarily, since $\rho_{f}$ is obtained by applying local quantum channels on a gapped topological order, no power-law correlation can be generated, i.e., $\rho_{f}$ exhibits short-range correlation for all local operators, which is a prerequisite for any topological order. Moreover, as we have seen, the subleading term in the entanglement negativity $\varepsilon_{A}(\rho_{f})$ is always $O(1)$ for $\rho_{f}$. In this sense, $\rho_{f}$ also retains certain essential properties of gapped topological order. Therefore, $\rho_{f}$ indeed represents a new type of topological order that is only possible in mixed states. The similarities and differences between gapless spin liquid and intrinsic mixed-state quantum TO are summarized in Table 1.
Gapless spin liquid Intrinsic mixed-state quantum TO
@@
Table 1: Comparison between the gapless spin liquid and the intrinsic mixed-state quantum TO ($\rho_{f}$). See Section III for properties listed in the last two rows

### v3

# Intrinsic Mixed-state Topological Order

@@ Abstract
Decoherence is a major obstacle to the preparation of topological order in noisy intermediate-scale quantum devices. Here, we show that decoherence can also give rise to new types of topological order. Specifically, we construct concrete examples by proliferating fermionic anyons in the toric code via local quantum channels. The resulting mixed states retain long-range entanglement, which manifests in the nonzero topological entanglement negativity, though the topological quantum memory is destroyed by decoherence. By comparison to the gapless spin liquid in pure states, we show that the identified states represent a novel intrinsic mixed-state topological order, which has no counterpart in pure states. Through the lens of quantum anomalies of 1-form symmetries, we then provide general constructions of intrinsic mixed-state topological order, and reveal the existence of non-bosonic deconfined anyons as another key feature of these novel phases. The extended meaning and characterization of deconfined excitations and their statistics in mixed states are clarified. Moreover, when these deconfined anyons have nontrivial braiding statistics, we prove that the mixed states cannot be prepared via finite-depth local quantum channels from any bipartite separable states. We further demonstrate our construction using the decohered Kitaev honeycomb model and the decohered double semion model. In the latter case, a surprising scenario arises where decoherence gives rise to additional types of deconfined anyons.
@@ I Introduction
As long-range entangled (LRE) quantum matter, topologically ordered phases have attracted extensive attention in the past few decades WEN (1990); Chen et al. (2010); Wen (2017); Savary and Balents (2016); Zhou et al. (2017); Sachdev (2018). Recently, there is a growing number of theoretical proposals Aguado et al. (2008); Verresen et al. (2021); Piroli et al. (2021); Tantivasadakarn et al. (2024, 2023a, 2023b); Lee et al. (2022a); Bravyi et al. (2022); Lu et al. (2023) as well as experimental evidence Semeghini et al. (2021); Satzinger et al. (2021); Andersen et al. (2023); Xu et al. (2023); Iqbal et al. (2023a, b) showing that topological order (TO) can be prepared in current many-body quantum simulation platforms, such as superconducting qubit arrays, Rydberg atom arrays, and trapped ion systems, etc. A key feature of these noisy intermediate-scale quantum (NISQ) devices is the inevitable presence of decoherence, which renders the quantum state a mixed state Preskill (2018). Substantial progress has been made in diagnosing nontrivial topological phases subject to decoherence Dennis et al. (2002); Bombin et al. (2012); McGinley and Cooper (2020); Deng et al. (2021); Wang et al. (2021); de Groot et al. (2022); Lee et al. (2022b); Zhang et al. (2022); Fan et al. (2024); Bao et al. (2023); Lee et al. (2023); Ma and Wang (2023); Ma et al. (2023); Bardyn et al. (2018); Mao et al. (2023); Su et al. (2023); Wang et al. (2023); Liu and Lieu (2024). Particularly, the decoherence-induced breakdown of topological quantum memory in the toric code model is investigated, and is related to the transition in the mixed-state topological order Fan et al. (2024); Lee et al. (2023). The topological order therein is inherited from the pure-state counterpart, which is resistant to modest decoherence. Above certain critical error rate, the long-range entanglement is destroyed.
@@
We are interested in the following question: Other than destroying the pure-state topological order, can decoherence give rise to novel types of topological order that are intrinsically mixed? The possibility of such an intriguing scenario arises from new mechanisms of anyon proliferation provided by decoherence, distinct from anyon condensation in pure states Bais and Slingerland (2009); Kong (2014); Burnell (2018). As a starting point, we explore such possibility in the context of $\mathbb{Z}_{2}$ (toric-code) topological order, which comprises three types of anyon excitations, $e,m,$ and $f=e\times m$, Anderson (1973); Read and Chakraborty (1989); Read and Sachdev (1991); Wen (1991); Anderson (1987); Senthil and Fisher (2000); Moessner and Sondhi (2001); Kitaev (2003). In pure states, either $e$ or $m$ (both being self-bosons) can condense, leading to a topologically trivial Higgs/confined phase Fradkin and Shenker (1979). Correspondingly, proliferation of $e,m$ anyons induced by decoherence also destroys long-range entanglement. In contrast, $f$ anyons are self-fermions, and therefore they cannot condense in a pure state; instead, strong fluctuation of $f$ anyons typically leads to a gapless spin liquid, which remains LRE. This motivates us to study the fate of $\mathbb{Z}_{2}$ topological order when $f$ anyons proliferate under decoherence.
@@
Specifically, we study the behavior of the toric code model Kitaev (2003) under local quantum channels that solely create $f$ anyons. The topological quantum memory degrades to classical memory above certain decoherence threshold, aligning with previous studies. However, it turns out that the mixed state still possesses nontrivial quantum topological order even in the absence of quantum memory. To diagnose the mixed-state topological order, we employ the topological entanglement negativity (TEN) Lee and Vidal (2013); Castelnovo (2013); Wen et al. (2016a, b), which is a natural generalization of the topological entanglement entropy (TEE) Kitaev and Preskill (2006); Levin and Wen (2006). TEN has been utilized to probe topological order in thermal equilibrium Hart and Castelnovo (2018); Lu et al. (2020, 2022); Lu and Vijay (2022) or under shallow-depth noise channels Fan et al. (2024), effectively acting as a faithful indicator of topological quantum memory in those studied cases. However, we find that the TEN fails to reflect topological quantum memory in this scenario, and remains unchanged across the transition. Nevertheless, the nonzero TEN still points to the persistence of long-range entanglement, a hallmark of topological order. Moreover, the absence of topological quantum memory indicates that the identified topological order has no pure-state counterpart and, therefore, is termed “intrinsic mixed-state topological order” here.
@@
We then provide further understanding of this peculiar result from the perspective of quantum anomalies. Crucially, the noisy channel proliferating $f$ anyons preserves an anomalous $1$-form symmetry generated by $f$ anyons Gaiotto et al. (2015); Wen (2019); McGreevy (2023). Through the anomalous $1$-form symmetry, we show that the intrinsic mixed-state TO in the decohered toric code supports deconfined fermionic anyons, with a detailed explanation of the extended meaning of deconfinement and fermionic statistics in mixed states. We generally prove that mixed states with anomalous $1$-form symmetries must be LRE. Moreover, we show that when the anomalous $1$-form symmetry is generated by anyons with nontrivial braiding statistics, any two complementary parts of the mixed state are LRE (for arbitrary bipartition), generalizing the conjecture “mixed-state anomaly $\Rightarrow$ multipartite non-separability” in Lessa et al. (2024), which focuses on 00-form symmetries. With this perspective, our construction can be generalized to get other intrinsic mixed-state TO, characterized by deconfined anyons with nontrivial statistics.
After introducing the general construction, we give two more examples of intrinsic mixed-state TO, the decohered Kitaev honeycomb model and the decohered double semion model. In the latter example, we find more surprising features in mixed-state TO. In addition to peeling off anyons in the original TO, decoherence can even give birth to new types of deconfined anyons, which leads to a non-modular anyon theory in our case.
@@
The rest of the paper is organized as follows. In Section II we construct an intrinsic mixed-state TO by proliferating $f$ anyons in the toric code, and reveal its properties through investigations of information quantities including the coherent information and the TEN. We also discuss similarities and disparities to the gapless spin liquid phase in pure states. In Section III we discuss general aspects of intrinsic mixed-state from the perspective of 1-form symmetry anomalies and deconfined excitations. In Section IV we generalize our construction of mixed-state quantum TO to decohered Kitaev honeycomb model and decohered double semion model. We conclude with a discussion in Section V. We note that Secs. III and IV.2 have been added after the initial version of our work appeared on arXiv. Meanwhile, Refs. Sohal and Prem (2024); Ellison and Cheng (2024) appeared, which study the mixed-state TO from the perspective of anomalous 1-form symmetries. The discussion regarding the nonmodular anyon theory in Sec. IV.2 is inspired by these works.
II Decohered Toric code as intrinsic mixed-state TO
@@ II.1 The model
$\mathcal{N}^{x}$ $=\prod_{i}\mathcal{N}_{i}^{x},\quad\mathcal{N}_{i}^{x}[\cdot]\equiv(1-p_{x})\cdot+p_{x}X_{i}\cdot X_{i},$ (1)
$\mathcal{N}^{z}$ $=\prod_{i}\mathcal{N}_{i}^{z},\quad\mathcal{N}_{i}^{z}[\cdot]\equiv(1-p_{z})\cdot+p_{z}Z_{i}\cdot Z_{i}.$
@@
where $\bm{\delta}=(\frac{1}{2},-\frac{1}{2})$ (the lattice constant is taken to be 1) and $0<p_{f}<\frac{1}{2}$ is the error rate. In this way the ground state $\rho_{0}$ is turned into a mixed state, $\rho_{f}=\mathcal{N}^{f}[\rho_{0}]$. As depicted in Fig. 1(b), $\mathcal{N}^{f}$ exclusively creates $f$ anyons. In contrast, certain other types of errors, like the Pauli-Y errors, locally create pairs of $f$ anyons as well, but globally they also produce $e$ and $m$ anyons, resulting in completely different outcomes. In the following sections, we demonstrate that this simple model surprisingly realizes an exotic intrinsic mixed-state topological order through analytical exact investigations of its topological memory and topological entanglement negativity.
@@ II.3 Topological entanglement negativity
Based on the above analysis, it may seem that $\rho_{f}$ closely resembles $\rho_{e}$ and $\rho_{m}$. However, surprisingly, we demonstrate below that even when the quantum memory breaks down for $p_{f}>p_{c}$, $\rho_{f}$ retains LRE with a nonzero TEN, indicating the emergence of a distinct topological order.
@@
$G_{x}\equiv\langle\{A_{v}\}\rangle,\quad G_{z}\equiv\langle\{B_{p}\}\rangle.$ (6)
@@
The dependence of TEN on the boundary size for the second type of bipartition seems a bit puzzling. However, this phenomenon also shows up in some ground-state topological order. Namely, the TEE/TEN also exhibits similar even/odd dependence on system size for $\mathbb{Z}_{2}$ topological order enriched by translation symmetry through weak symmetry breaking, with typical examples including the Wen-plaquette model and the Abelian phase of Kitaev honeycomb model Wen (2003); Kitaev (2006). In these cases, it is well known that the ground state degeneracy on a torus also exhibits similar dependence on the system size. Although there is little discussion in the literature about how weak symmetry breaking affects entanglement properties, it is straightforward to check that the subleading term of bipartite entanglement entropy on a cylinder also depends on the parity of the boundary size in these models. Based on this observation, we establish a connection between the entanglement properties of $\rho_{f}$ and ground-state $\mathbb{Z}_{2}$ topological order enriched by translation symmetry in Appendix A.4.
@@
II.4 Robustness of the intrinsic mixed-state TO
Although in our construction we need to use specific $2$-qubit channels that look a bit unconventional, $\rho_{f}$ represents a new type of topologically ordered phase, instead of a fine-tuned exception. We can consider the case when single-qubit phase errors are also present: ${\rho}_{f,e}=\mathcal{N}^{z}[\rho_{f}]$, with error rate $p_{z}$. By mapping to two decoupled RBIMs, we obtain the phase diagram in Fig. 2. For small $p_{z}$, ${\rho}_{f,e}$ stays in the same phase as $\rho_{f}$, while for $p_{z}>p_{c}\approx 0.109$, the state undergoes another transition to the trivial phase, with no memory and zero TEN. See Appendix A.5 for details.
Figure 2: The phase diagram of the toric code model subjected to both two-qubit errors and single-qubit phase errors. Properties of each phase (including topological memory and TEN) and the corresponding phases in the two RBIMs (with Ising variables $\sigma$,$\tau$, respectively) are indicated in the phase diagram.
@@ II.5 Comparison to anyon condensation and gapless spin liquid
The distinction becomes much more significant for $f$ anyons. As fermions, they cannot condense in pure states. Then what happens when we turn on fluctuations of $f$ anyons ($h_{zx}\neq 0$)? Surprisingly, at $h_{z}=h_{x}=0$, this model can be exactly solved via fermionization Chen et al. (2018). As we show in Appendix B, when $h_{xz}$ is sufficiently large ($h_{x,z}=0$), corresponding to large fluctuations of $f$ particles, the system enters a gapless spin liquid phase, where $f$ particles form a $p-$wave superconductor with conic dispersion, similar to the gapless Kitaev spin liquid Kitaev (2006). The intrinsic mixed-state TO proposed in this paper has many common features with the gapless spin liquid. First, they are both obtained by proliferating $f$ anyons in the $\mathbb{Z}_{2}$ TO, and, as a result, they both have no quantum memory. For the gapless spin liquid, it is due to the absence of a spectrum gap, so there is no well-defined topological degeneracy. Second, despite the lack of quantum memory, they are both LRE, and are thus nontrivial phases of matter. In the next section we will uncover a deeper reason of their LRE nature from the perspective of anomalies. However, there are also notable differences between $\rho_{f}$ and the gapless spin liquid phase. Gapless phases typically exhibit critical behaviors, including algebraically decaying correlation functions and subleading logarithmic corner contributions to entanglement entropy/negativity Fradkin and Moore (2006); Casini et al. (2009); Bueno et al. (2015). Contrarily, since $\rho_{f}$ is obtained by applying local quantum channels on a gapped topological order, no power-law correlation can be generated, i.e., $\rho_{f}$ exhibits short-range correlation for all local operators, which is a prerequisite for any topological order. Moreover, as we have seen, the subleading term in the entanglement negativity $\varepsilon_{A}(\rho_{f})$ is always $O(1)$ for $\rho_{f}$. In this sense, $\rho_{f}$ also retains certain essential properties of gapped topological order. Therefore, $\rho_{f}$ indeed represents a new type of topological order that is only possible in mixed states. The similarities and differences between gapless spin liquid and intrinsic mixed-state TO are summarized in Table 1.
Gapless spin liquid Intrinsic mixed-state TO
@@
Table 1: Comparison between the gapless spin liquid and the intrinsic mixed-state TO ($\rho_{f}$). See Section III for properties listed in the last two rows